MONEY AND INVESTMENT
4) Maria has twice as much money invested at 5% simple annual interest as she does at 4%. If her yearly income from these two investments is Php 350, how much does she have invested at each rate?


Thank you in advance! Hoping for your immediate reply.

Answer by Theo(13342)

x = amount of money invested at 4%
2*x = amount of money invested at 5%

yearly income = 350.

.04*x + .05*2*x = 350

factor out the x to get:

x * (.04 + .05*2) = 350

simplify to get:

x * (.14) = 350

divide both sides by .14 to get:

x = 350 / .14 = 2500

that makes 2x = 5000

she invested 2500 at 4% and 5000 at 5%

.04 * 2500 + .05 * 5000 = 100 + 250 = 350

Answer is confirmed.

she invested 2500 at 4% and 5000 at 5%.